# The Open Journal of Astrophysics

### v2

@@ 3 The IAEmu Model
Figure 2: Model Pipeline. The HOD input model parameters are normalized before entering the 7-layer deep fully-connected embedding network. The embedding network expands the dimensionality of the input before a bottleneck latent space that transitions to the decoder stage, which features seven 1D convolutional layers which learn the individual local correlations present in the output correlation functions, $\log\xi$, $\omega$, and $\eta$. Both the embedding network and decoder feature residual connections, visualized as dotted arrows, to aid the convergence of IAEmu during training. IAEmu is trained using the $\beta$-NLL loss (Seitzer et al., 2022) with a 100 epoch warm-up period corresponding to mean-squared-error optimization before re-introducing aleatoric uncertainties into the optimization. The outputted correlation functions are then re-scaled back to their original values. A detailed description of the model training procedure is shown in Appendix C. N-body simulation visualization in the right panel is from (Perraudin et al., 2019).
@@ 3.1 Model Architecture
Our objective is to construct a neural network (NN) that replicates the mapping between HOD simulations and correlation function estimators. Specifically, the NN will take a 7-dimensional input vector of galaxy HOD and IA parameters, as described in Section 2 and illustrated in Figure 2, and predict the correlation functions $\xi(r)$, $\omega(r)$, and $\eta(r)$ across 20 bins. We represent each correlation function by a vector recording the value for 20 evenly spaced values of $r$. Additionally, the model directly outputs predictions of the aleatoric uncertainties $\sigma^{\text{aleo}}$ of the correlation amplitudes. This allows us to capture the stochastic nature of HOD modeling through a mean-variance estimation (MVE) training procedure (Nix & Weigend, 1994). Separately, we use Monte Carlo dropout to track the epistemic uncertainties $\sigma^{\text{epi}}$ inherent in the NN model. These can arise from limited training data or architecture misspecification. Both types of uncertainties are useful for analyzing $\omega(r)$ and $\eta(r)$ performance, which are inherently noisier statistics due to the significant effects of galaxy shape noise in correlations (Bernstein & Jarvis, 2002). Mathematically, the task mapping is a function
@@
We implement $f_{\phi}$ as an NN called IAEmu using PyTorch (Paszke et al., 2019). The IAEmu architecture includes a fully connected embedding network and three 1D convolutional NN decoder heads, trained using a multitask learning approach as shown in Figure 2.
The embedding network contains five fully connected linear layers, each followed by batch normalization and LeakyReLU activation (Xu et al., 2015). There are also residual connections in the third and fourth layer, allowing for improved information flow and gradient stability across the network (He et al., 2015). The embedding network increases the size of the input vector $X\in\mathbb{R}^{7}$ layer-by-layer to a $256$-dimensional latent feature, which is then mapped through a final bottleneck layer to a 128-dimensional latent vector $v$. To mitigate overfitting, we implement dropout (Srivastava et al., 2014). As described later, we also use dropout to isolate the epistemic uncertainty associated with the model’s parameters through the Monte Carlo dropout technique (Gal & Ghahramani, 2016).
The decoders each contain seven 1D convolutional decoder layers. Each decoder first takes the output of the embedding network, a feature vector $v$ of size 128, and maps it into an expanded feature space. This expanded feature vector is then reshaped to create a multi-channel 1D feature map, enabling the decoder to utilize 1D convolution to spatially transform the latent representation. Each layer has batch normalization, LeakyReLU activation, and dropout. Residual connections are introduced by adding the output of the second convolutional layer to the output of the third layer and by adding the output of the fifth layer to the output of the sixth layer. Each decoder gradually downsamples the latent representation $v$ and finally outputs a 2-channel 1D signal as a tensor of shape $2\times 20$ where the 2 channels represent the correlation amplitudes and variances respectively of the correlation function. To ensure variances are strictly positive, they are passed through a softplus activation in the output layer.
The IAEmu design serves a dual purpose, it facilitates vector-to-sequence conversion through the convolution of encoded representations from the embedding network and, within our multitask framework, enables separate forward paths to isolate features unique to each individual correlation estimator.
@@ 3.2 Training
$x^{\prime}=\frac{x-\mu_{x}}{\sigma_{x}}\;,$ (5)
where $\mu_{x}$ and $\sigma_{x}$ are the mean and standard deviation of the respective feature across the entire training dataset. This is an affine transformation and is thus easily invertible.
@@
To predict the mean and variance of the values of the correlation function, we use the $\beta$-NLL loss from (Seitzer et al., 2022), which is defined as
@@
This is similar to Gaussian-NLL loss (Nix & Weigend, 1994), defined
@@
where $X$ denotes the input data vector, $\hat{\mu}(X)$ and $\hat{\sigma}^{2}(X)$ the model predictions at an individual bin, and $Y$ the ground truth label. The numerator of the second term in Equation 7 is the typical mean-squared-error loss, used when the model only outputs a point estimate approximating the mean of the distribution. One drawback of the Gaussian NLL loss is that the model can become stuck in local minima in the loss landscape during training. This results in a prediction with an incorrect mean and high variance. However, by adjusting $\beta$ appropriately, this risk can be reduced. The utility of the $\beta$-NLL loss can be seen in the gradients:
$\nabla_{\hat{\mu}}\mathcal{L}_{\beta\text{-NLL}}(\theta)$ $=\mathbb{E}_{X,Y}\left[\frac{\hat{\mu}(X)-Y}{\hat{\sigma}^{(2-2\beta)}(X)}\right]$
$\nabla_{\hat{\sigma}^{2}}\mathcal{L}_{\beta\text{-NLL}}(\theta)$ $=\mathbb{E}_{X,Y}\left[\frac{\hat{\sigma}^{2}(X)-(Y-\hat{\mu}(X))^{2}}{2\hat{\sigma}^{(4-2\beta)}(X)}\right]\;.$
The $\beta$ parameter allows one to interpolate between Gaussian-NLL in the limit that $\beta\rightarrow 0$, and standard MSE in the limit that $\beta\rightarrow 1$. This loss has the benefit of allowing one to encode the strength of the mean prediction to the loss, to discourage local minima with poor mean predictions and large variances. It was empirically found in Seitzer et al. (2022) that a value of $\beta=0.5$ generally performs best. However, we explore different values of $\beta$ and introduce a warm-up period of $\ell^{\prime}$ epochs to enable individualized training for each correlation. The total loss function during training at epoch $\ell$ is:
$\mathcal{L}(\ell;\theta)=\begin{aligned} &\begin{cases}\begin{aligned} &\mathcal{L}_{\beta\text{-NLL}}^{\xi}(\theta,1.0)+\mathcal{L}_{\beta\text{-NLL}}^{\omega}(\theta,1.0)\\ &+\mathcal{L}_{\beta\text{-NLL}}^{\eta}(\theta,1.0)\end{aligned}&\text{if }\ell<\ell^{\prime}\\[5.69054pt] \begin{aligned} &\mathcal{L}_{\beta\text{-NLL}}^{\xi}(\theta,0.9)+\mathcal{L}_{\beta\text{-NLL}}^{\omega}(\theta,0.5)\\ &+\mathcal{L}_{\beta\text{-NLL}}^{\eta}(\theta,0.5)\end{aligned}&\text{if }\ell\geq\ell^{\prime}\end{cases}\end{aligned}$ (8)
@@ 4 Results
We analyze the performance of IAEmu, first on the held-out (in-distribution) test set, and further on a set of IA observations from the IllustrisTNG suite of hydrodynamical simulations. IAEmu achieves high average accuracy for both galaxy position-position and position-orientation statistics, and demonstrates robustness to shape noise in the orientation-orientation statistics without signs of overfitting. We show that IAEmu’s performance on $\eta$ is more difficult to quantify due to the high stochasticity of the correlation function, even after averaging over multiple realizations of the data. We lastly show that when fitting alignment parameters to IA correlations from IllustrisTNG, IAEmu has better than $0.4\sigma$ agreement with halotools-IA.
@@ 4.1 Performance
Accuracy. We evaluate the model on the 20% in-distribution but held-out test set, as summarized in Figure 3. All test-set predictions are mean predictions averaged over 50 forward passes (i.e., predictions with Monte Carlo Dropout) of IAEmu, so that an epistemic uncertainty on predictions can be retrieved. In reporting metrics, we exclude outlier examples in the $\omega$ and $\eta$ correlations where IAEmu shows a strong Spearman correlation coefficient ($>0.5$) with the ground truth or its predicted amplitude is within $1\sigma$ of the true uncertainty of the data for the majority of the bins, but the fractional error exceeds a factor of $100$ and $450$, respectively. These extreme values would thus be artifacts from the small amplitude and inherent variability of the ground truth data, rather than actual model inaccuracies. These thresholds were chosen as they remove only $\lesssim 1\%$ of test-set predictions, while stabilizing the mean fractional error on the test set. Even with this mitigation, many instances remain in the test set where the performance of IAEmu is visually suitable, but features large fractional error due to these numerical artifacts.
@@
For the position-position ($\xi$) correlation, the mean fractional error per bin (top panel) reaches a maximum of $10\%$, with IAEmu achieving an average error of $3.2\%$ for $\xi$. The $\xi$ performance is biased high at large $r$, where the $\xi$ correlation amplitudes are small ($||\xi||\ll 1$) and approach zero. This bias may arise in part from the standardization for training, which can disproportionately emphasize regions with larger amplitudes or compress the dynamic range at small values, leading to systematic bias. Additionally, as IAEmu naturally predicts $\log\xi$, small residuals near zero can appear large when transformed back to linear space.
@@
For $\omega$, the accuracy drops at the onset of the 2-halo regime, with an average model error across all bins of $4.9\%$ and a similar maximum of $\sim 10\%$. We also find that the median fractional error across all bins is less than $10\%$ for $66\%$ of test-set predictions. This is approaching the accuracy for IA modeling likely required for Stage IV surveys (Paopiamsap et al., 2024). The mean fractional error for orientation-orientation ($\eta$) is significantly higher, averaging $54\%$. However, it is important to note that the ground truth $\omega$ and $\eta$ correlations – even after averaging over 10 realizations of the dataset – are generally noisy and can often fluctuate between positive and negative values. Fractional error can be an inappropriate measure in this case, as it becomes ill-defined when the amplitude of the data is close to zero or changes sign, leading to misleadingly large error values.
@@
With this in mind, we show in the middle panel of Figure 3 the median residual in units of the dataset’s true aleatoric uncertainty $\hat{\sigma}$. From this metric, it is observed that despite the large fractional error in $\eta$, the predictions of IAEmu remain strictly within $1\sigma$ of the ground truth correlations across all bins. This trend also holds for $\omega$. For $\xi$, the residual is computed in log space in the 2-halo regime to more consistently represent the bias with how IAEmu was trained, and to avoid exaggerated deviations caused by exponentiating small correlation values. Despite the large stochasticity of $\omega$ and $\eta$, this indicates that IAEmu has learned to capture the mean behavior and not overfit to the noise fluctuations in these correlations. This provides the added benefit of capturing the “cosmic mean” of the correlations directly with IAEmu, which would otherwise require running multiple realizations of the underlying HOD. This can also be frequently seen in example IAEmu predictions for $\omega$ and $\eta$ as seen in Appendix D.
@@
where $y_{i}$ represents the ground truth value, and $\hat{y}_{i}$ denotes the corresponding prediction by IAEmu. This metric provides an indication of prediction accuracy, but can be sensitive to outliers. To quantify relative percentage error, we use the SMAPE, which is defined
@@
where $\epsilon=10^{-8}$ is introduced to prevent division by zero. The SMAPE is generally more robust to outliers compared to the NRMSE, but it tends to be more sensitive to small values. These three metrics are selected due to their scale-invariant properties, which are essential for comparing IAEmu’s performance across the varying scales of $\xi$, $\omega$, and $\eta$. An SCC value of 1 indicates a perfect correlation between IAEmu and the ground truth data, while lower values of NRMSE and SMAPE reflect better predictive performance. Together, these metrics provide a comprehensive assessment of IAEmu’s performance.
@@
For $\xi$, we find a SCC value of $0.99$ when averaged across all bins, and a value of $0.98$ for $\omega$ as seen in Figure 3. This indicates a very strong correlation between the IAEmu predictions and the underlying data. For $\eta$, the average SCC across all bins is $0.55$, with the $\text{SCC}\approx 0.75$ at low $r$, but is around 0.5 after entering the 2-halo regime, which still reflects a moderate correlation between the data and model. At larger $r$, the SCC decreases, indicating a weak correlation. It is important to note that the SCC can be strongly affected by stochasticity and the low amplitude of the data, particularly when the amplitudes approach zero, as is the case frequently for $\eta(r)$.
@@
For $\xi$ and $\omega$, the NRMSE averaged across all bins is $0.19$, as shown in the bottom panel of Figure 3. The corresponding SMAPE values are $\text{SMAPE}(\xi)=0.10$ and $\text{SMAPE}(\omega)=0.30$. The relatively low NRMSE values indicate that, on average, the predictions closely follow the ground truth across the full range of data. However, the higher SMAPE for $\omega$ compared to $\xi$ suggests that the relative error is more pronounced for $\omega$, potentially due to the generally smaller correlation amplitudes in $\omega$. This implies that while the absolute prediction error remains comparable, the percentage error is exacerbated by the lower magnitude of the true values in $\omega$. A similar trend is observed for $\eta$, where both the NRMSE ($0.69$) and SMAPE ($1.11$) are significantly larger. These higher values indicate that IAEmu’s predictions for $\eta$ exhibit larger absolute and relative deviations from the ground truth. This could be attributed to a decrease in performance, increased variability, or a broader dynamic range in $\eta$, which naturally poses greater challenges for accurate predictions.
@@
Limitations. IAEmu’s predictions for $\eta$ are less accurate compared to $\xi$ and $\omega$, which perform well across metrics considered. While IAEmu successfully captures the correct scaling of $\eta$ across all bins, its accuracy for $\omega$ and $\eta$ is primarily limited by the stochastic nature of these correlations, even when trained on multiple realizations and evaluated on their means. As demonstrated by examples in Appendix D, the averaged ground truth correlations still exhibit fluctuations that are indicative of noise due to the relatively small volume considered for the simulations. This hinders the evaluation of IAEmu’s performance as well as training; however, as also demonstrated in the middle panel of Figure 3, IAEmu reliably captures the underlying mean behavior despite the presence of noise. The bias in at large $r$ for $\xi$ can likely be attributed to the use of $\log$ and the standardization procedure for training. In a multi-task framework, standardization can potentially be avoided by using trainable loss coefficients (Kendall et al., 2018).

### v4

@@ 3 The IAEmu Model
Figure 2: Model Pipeline. The HOD input model parameters are normalized before entering the 7-layer deep multilayer perceptron (MLP) embedding network. The embedding network expands the dimensionality of the input before a bottleneck latent space that transitions to the decoder stage, which features seven 1D convolutional layers which learn the individual local correlations present in the output correlation functions, $\log\xi$, $\omega$, and $\eta$. Both the embedding network and decoder feature residual connections to aid the convergence of IAEmu during training. IAEmu is trained using the $\beta$-NLL loss (seitzer2022pitfallsheteroscedasticuncertaintyestimation) with a 100 epoch warm-up period corresponding to mean-squared-error optimization before re-introducing aleatoric uncertainties into the optimization. The generated correlation functions are then re-scaled back to their original values. A detailed description of the model training procedure is shown in Appendix C. NN-body simulation visualization in the right panel is from (nbody).
@@ 3.1 Model Architecture
Our objective is to construct a neural network (NN) that replicates the mapping between HOD simulations and correlation function estimators. Specifically, the NN will take a 7-dimensional input vector of galaxy HOD and IA parameters, as described in Section 2 and illustrated in Figure 2, and predict the correlation functions $\xi(r)$, $\omega(r)$, and $\eta(r)$ across 20 $r$ bins. We represent each correlation function by a vector recording the value for 20 evenly spaced values of $r$. Additionally, the model directly outputs predictions of the aleatoric uncertainties $\sigma^{\text{aleo}}$ on the correlation amplitudes. This allows us to capture the stochastic nature of HOD modeling through a mean-variance estimation (MVE) training procedure (MVE). Separately, we use Monte Carlo dropout to track the epistemic uncertainties $\sigma^{\text{epi}}$ inherent in the NN model. These can arise from limited training data or architecture misspecification. Both types of uncertainties are useful for analyzing $\omega(r)$ and $\eta(r)$ performance, which are inherently noisier statistics due to the significant effects of galaxy shape noise in correlations (Bernstein_2002). Mathematically, the task mapping is a function
@@
We implement $f_{\phi}$ as an NN called IAEmu using PyTorch (paszke2019pytorchimperativestylehighperformance). The IAEmu architecture includes a fully connected embedding network and three 1D convolutional NN decoder heads, trained using a multitask learning approach as shown in Figure 2.
The embedding network contains five fully connected linear layers, each followed by batch normalization and LeakyReLU activation (xu2015empirical). Residual connections link the second and third layers using a linear projection to match dimensions, and the third and fourth layers by directly adding the layer outputs, which improves information flow and gradient stability (he2015deepresiduallearningimage). The embedding network increases the size of the input vector $X\in\mathbb{R}^{7}$ layer-by-layer to a $256$-dimensional latent feature, which is then mapped through a final bottleneck layer to a 128-dimensional latent vector $v$. To mitigate overfitting, we incorporate dropout (JMLR:v15:srivastava14a) into the IAEmu architecture during training. Additionally, as detailed later, we leverage dropout to estimate the epistemic uncertainty associated with the model’s parameters using the Monte Carlo dropout technique (gal2016dropout).
The decoders each contain seven 1D convolutional decoder layers. Each decoder first takes the output of the embedding network, a feature vector $v$ of size 128, and maps it into an expanded feature space. This expanded feature vector is then reshaped to create a multichannel 1D feature map, enabling the decoder to utilize 1D convolution to spatially transform the latent representation. Each layer has batch normalization, LeakyReLU activation, and dropout. Residual connections are introduced by adding the output of the second convolutional layer to the output of the third layer and by adding the output of the fifth layer to the output of the sixth layer. Each decoder gradually downsamples the latent representation $v$ and finally outputs a 2-channel 1D signal as a tensor of shape $2\times 20$, where the 2 channels represent the correlation amplitudes and variances of the correlation function, respectively. To ensure variances are strictly positive, they are passed through a softplus activation in the output layer.
The IAEmu design serves a dual purpose: it facilitates vector-to-sequence conversion through the convolution of encoded representations from the embedding network and, within our multitask framework, enables separate forward paths to isolate features unique to each individual correlation estimator.
@@ 3.2 Training
$z=\frac{x-\mu_{x}}{\sigma_{x}}\;,$ (5)
where $\mu_{x}$ and $\sigma_{x}$ are the mean and standard deviation of the respective feature across the entire training dataset. This is known as $z$-score standardization; it is an affine transformation and is thus easily invertible.
@@
To predict the mean and variance of the values of the correlation function, we use the $\beta$-NLL loss from (seitzer2022pitfallsheteroscedasticuncertaintyestimation), which is defined as
@@
This is similar to Gaussian-NLL loss (MVE), defined
@@
where $X$ denotes the input data vector, $\hat{\mu}(X)$ and $\hat{\sigma}^{2}(X)$ the model predictions at an individual bin, $Y$ the ground truth label, and $C$ a normalization constant. The numerator of the second term in Equation 7 is the typical mean-squared-error loss, used when the model only outputs a point estimate approximating the mean of the distribution. One drawback of the Gaussian NLL loss is that the model can become stuck in local minima in the loss landscape during training. This results in a prediction with an incorrect mean and high variance. However, by adjusting $\beta$ appropriately, this risk can be reduced. The utility of the $\beta$-NLL loss can be seen in the gradients:
$\nabla_{\hat{\mu}}\mathcal{L}_{\beta\text{-NLL}}(\theta)$ $=\mathbb{E}_{X,Y}\left[\frac{\hat{\mu}(X)-Y}{\hat{\sigma}^{(2-2\beta)}(X)}\right]$ (8)
$\nabla_{\hat{\sigma}^{2}}\mathcal{L}_{\beta\text{-NLL}}(\theta)$ $=\mathbb{E}_{X,Y}\left[\frac{\hat{\sigma}^{2}(X)-(Y-\hat{\mu}(X))^{2}}{2\hat{\sigma}^{(4-2\beta)}(X)}\right]\;.$ (9)
The $\beta$ parameter allows one to interpolate between Gaussian-NLL in the limit that $\beta\rightarrow 0$, and standard MSE in the limit that $\beta\rightarrow 1$. This loss has the benefit of allowing one to encode the contribution of the mean prediction to the loss, to discourage local minima with poor mean predictions and large variances. It was empirically found in seitzer2022pitfallsheteroscedasticuncertaintyestimation that a value of $\beta=0.5$ generally performs best. However, we explore different values of $\beta$ and introduce a warm-up period of $\ell^{\prime}$ epochs to enable individualized training for each correlation. The total loss function during training at epoch $\ell$ is:
[…]
@@ 4 Results
We analyze the performance of IAEmu, first on the held-out (in-distribution) test set, and further on a set of IA observations from the tng300 suite of hydrodynamical simulations. IAEmu achieves high average accuracy for both galaxy position-position and position-orientation statistics, and demonstrates robustness to shape noise in the orientation-orientation statistics without signs of overfitting. We show that IAEmu’s performance on $\eta$ is more difficult to quantify due to the high stochasticity of the correlation function, even after averaging over multiple realizations of the data. We lastly show that when fitting alignment parameters to IA correlations from tng300, IAEmu has better than $0.4\sigma$ agreement with halotools-IA.
@@ 4.1 Performance
Accuracy. We evaluate the model on the 20% in-distribution but held-out test set, as summarized in Figure 3. All test-set predictions are mean predictions averaged over 50 forward passes (i.e., predictions with Monte Carlo Dropout) of IAEmu, so that an epistemic uncertainty on predictions can be retrieved. Reported metrics are evaluated on the correlations in their original domain; they are not computed in the standardized or log-transformed domain employed during IAEmu training.
@@
In reporting metrics, we exclude outlier examples in the $\omega$ and $\eta$ correlations where IAEmu shows a strong Spearman correlation coefficient (>0.5>0.5) with the ground truth or its predicted amplitude is within $1\sigma$ of the true uncertainty of the data for the majority of the bins, but the fractional error exceeds a factor of $100$ and $450$, respectively. These extreme values arise from the lack of a high-signal-to-noise ground truth and occur when the true correlation amplitude is small and changes sign, conditions under which the fractional error can become arbitrarily large even when predictions accurately follow the underlying signal. They are therefore artifacts of the small amplitude and noise of the data, rather than indicators of genuine model inaccuracies. The thresholds were chosen empirically to exclude only these pathological cases, affecting fewer than $\lesssim 1\%$ of test-set predictions. Including these outliers does not impact the overall assessment of performance but significantly inflates the mean fractional error, introducing numerical instability. Notably, the median fractional error remains stable—within approximately $0.1\%$—for both $\omega$ and $\eta$, irrespective of whether these extreme cases are retained. Even with this mitigation, many instances remain in the test set where the performance of IAEmu is visually suitable, but features large fractional error due to these numerical artifacts.
@@
For the position-position ($\xi$) correlation, the mean fractional error per bin (top panel) reaches a maximum of $10\%$, with IAEmu achieving an average error of $3.2\%$ for $\xi$. The $\xi$ performance is biased high at large $r$, where the $\xi$ correlation amplitudes are small ($\xi||\ll 1$) and approach zero. This bias may arise in part from the standardization for training, which can disproportionately emphasize regions with larger amplitudes or compress the dynamic range at small values, leading to systematic bias. Additionally, as IAEmu naturally predicts $\log\xi$, small residuals near zero can appear large when transformed back to linear space.
@@
For $\omega$, the accuracy drops at the onset of the 2-halo regime, with an average model error across all bins of $4.9\%$ and a similar maximum of $\sim 10\%$. We also find that the median fractional error across all bins is less than $10\%$ for $66\%$ of test-set predictions. This is approaching the accuracy for IA modeling likely required for Stage IV surveys (Paopiamsap_2024). The mean fractional error for orientation-orientation ($\eta$) is significantly higher, averaging $54\%$. However, it is important to note that the ground truth $\omega$ and $\eta$ correlations – even after averaging over 10 realizations of the dataset – are generally noisy and can often fluctuate between positive and negative values. Fractional error can thus be misleading in this case due to the absence of high-signal ground truth values for comparison, and due to the correlation amplitudes being close to zero and frequently changing sign. We also studied the $r$-weighted mean fractional errors ((5.5%, 5.2%, and 67.4%) compared to their unweighted counterparts (3.2%, 4.9%, and 53.8%), which reveals a $\sim 2\%$ difference for $\xi$, a similar performance for $\omega$, and a larger difference for $\eta$. This is in line with the observed bias for $\xi$ at large $r$. In the case of $\eta$, we emphasize that fractional error for $\eta$ should not be interpreted in isolation as a gauge of model accuracy.
@@
With this in mind, we show in the middle panel of Figure 3 the median residual in units of the dataset’s true aleatoric uncertainty $\hat{\sigma}$. From this metric, it is observed that despite the large fractional error in $\eta$, the predictions of IAEmu remain strictly within $1\sigma$ of the ground truth correlations across all bins. This trend also holds for $\omega$. For $\xi$, the residual is computed in log space in the 2-halo regime to more consistently represent the bias with how IAEmu was trained, and to avoid exaggerated deviations caused by exponentiating small correlation values. Despite the large stochasticity of $\omega$ and $\eta$, this indicates that IAEmu has learned to capture the mean behavior and not overfit to the noise fluctuations in these correlations. This provides the added benefit of capturing the “cosmic mean” of the correlations directly with IAEmu, which would otherwise require running multiple realizations of the underlying HOD. This can also be frequently seen in example IAEmu predictions for $\omega$ and $\eta$ as shown in Appendix D.
@@
where $y_{i}$ represents the ground truth value, and $\hat{y}_{i}$ denotes the corresponding prediction by IAEmu. This metric provides an indication of prediction accuracy, but can be sensitive to outliers, due to dependence on squared error. To quantify relative percentage error, we use the SMAPE, which is defined
@@
where $\epsilon=10^{-8}$ is introduced to prevent division by zero. The SMAPE is generally more robust to outliers compared to the NRMSE, but it tends to be more sensitive to small values (i.e., small correlation amplitudes). These three metrics are selected due to their scale-invariant properties, which are essential for comparing IAEmu’s performance across the varying scales of $\xi$, $\omega$, and $\eta$. An SCC value of 1 indicates a perfect correlation between IAEmu and the ground truth data, while lower values of NRMSE and SMAPE reflect better predictive performance. The reported metrics are averaged over correlations in the test set. Together, these metrics provide a comprehensive assessment of IAEmu’s performance.
@@
For $\xi$, we find an SCC value of $0.99$ when averaged across all bins, and a value of $0.98$ for $\omega$ as seen in Figure 3. This indicates a very strong correlation between the IAEmu predictions and the underlying data. For $\eta$, the average SCC across all bins is $0.55$, with the $\text{SCC}\approx 0.75$ at low $r$, but it is around 0.5 after entering the 2-halo regime, which still reflects a moderate correlation between the data and model. At larger $r$, the SCC decreases, indicating a weak correlation. It is important to note that the SCC can be strongly affected by stochasticity and the low amplitude of the data, particularly when the amplitudes approach zero, as is the case frequently for $\eta(r)$.
@@
For $\xi$ and $\omega$, the NRMSE averaged across all bins is $0.19$, as shown in the bottom panel of Figure 3. The corresponding SMAPE values averaged across all bins are $\text{SMAPE}(\xi)=0.10$ and $\text{SMAPE}(\omega)=0.30$. The relatively low NRMSE values indicate that, on average, the predictions closely follow the ground truth across the full range of data. However, the higher SMAPE for $\omega$ compared to $\xi$ suggests that the relative error is more pronounced for $\omega$, potentially due to the generally smaller correlation amplitudes in $\omega$. This implies that while the absolute prediction error remains comparable, the percentage error is exacerbated by the lower magnitude of the true values in $\omega$. A similar trend is observed for $\eta$, where both the NRMSE ($0.69$) and SMAPE ($1.11$) are significantly larger. These higher values indicate that IAEmu’s predictions for $\eta$ exhibit larger absolute and relative deviations from the ground truth. This could be attributed to a decrease in performance, increased variability, or a broader dynamic range in $\eta$, which naturally poses greater challenges for accurate predictions.
@@
Limitations. IAEmu’s predictions for $\eta$ are less accurate compared to $\xi$ and $\omega$, which perform well across metrics considered. While IAEmu successfully captures the correct scaling of $\eta$ across all bins, its accuracy for $\omega$ and $\eta$ is primarily limited by the stochastic nature of these correlations, even when trained on multiple realizations and evaluated on their means. As demonstrated by examples in Appendix D, the averaged ground truth correlations still exhibit fluctuations that are indicative of noise due to the relatively small volume considered for the simulations. This hinders the evaluation of IAEmu’s performance as well as training; however, as also demonstrated in the middle panel of Figure 3, IAEmu reliably captures the underlying mean behavior despite the presence of noise. The bias at large $r$ for $\xi$ can likely be attributed to the use of $\log$ and the standardization procedure for training. In a multitask framework, standardization can potentially be avoided by using trainable loss coefficients (kendall2018multitasklearningusinguncertainty).